# MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation

### v1

@@ Abstract
With the rise of online dance‑video platforms and rapid advances in AI‑generated content (AIGC), music‑driven dance generation task has emerged as a compelling research direction. Despite substantial progress in related domains such as music-driven 3D dance generation, pose-driven image animation, and audio-driven talking-head synthesis, they cannot be directly adapted to this task. Moreover, the limited studies in this task still struggle to jointly achieve high-quality visual appearance and human motion. Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training (GFT) strategy, achieving state-of-the-art (SOTA) performance in 3D dance generation task; the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance in the pose-driven image animation task. To better benchmark this task, we curate a large-scale and diverse dataset, and design a motion–appearance evaluation protocol. Based on them, MACE-Dance also achieves the state-of-the-art (SOTA) performance. Project page: https://macedance.github.io/.
@@
Figure 1: Leveraging the synergistic collaboration among the cascaded experts, MACE-Dance can generate diverse dance videos that not only exhibit kinematically plausible and artistically expressive motion, but also maintain spatiotemporal coherent appearance.
1 Introduction
@@
Dance is a vital part of human culture. Moving to the beat and melody, dancers both convey emotion and narrative intent and showcase the power and beauty of human movement [37, 2, 1]. In the era of the internet, dance videos have become highly prominent on platforms such as YouTube and TikTok. In parallel, rapid advances [46, 48] in AI-generated content (AIGC) has created the technical preconditions for automating dance video creation, making it a timely and impactful research direction. Nevertheless, this task faces two key challenges [45]: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency.
@@
Recent progress in dance generation has focused primarily on 3D dance [37, 19, 17], with numerous strong methods emerging across model families—autoregressive [30, 47, 46], GAN-based [44, 35, 13], and diffusion-based [37, 19, 18]. Although 2D dance videos can be rendered from 3D motion, the resulting outputs typically miss realistic human–scene interactions and high-fidelity human textures, leading to a suboptimal visual appearance [45]. In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance. Traditionally, pose-driven image animation has achieved notable advances [36, 12, 5]. However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—remains manual [2]. Similarly, audio-driven talking head generation has also achieved significant breakthroughs [26, 27, 28]. However, its direct transfer to dance video generation remains challenging, as it primarily focuses on relatively simple upper-body gesture rather than the complex full-body motion required in dance [25, 24]. Research on music-driven dance video generation remains limited [4, 40], and there is still insufficient exploration of jointly achieving high-quality visual appearance and human motion.
@@
Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.   MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation . The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. (1) Motion Expert. Motion Expert adopts Diffusion Model with BiMamba-Transfomer hybrid architecture. The bidirectional Mamba [8] captures intra-modal local dependencies in music or dance, while the Transformer [39] models cross-modal global context. Owing to this architecture, the Motion Expert generates entire sequence in non-autoregressive manner during inference, not only improving generation efficiency, but also avoiding the exposure bias problem in autoregressive [47] and inpainting-based [37] methods. To enhance generation stability and accelerate inference, we employ guidance-free training (GFT [3]) instead of conventional classifier-free guidance (CFG [9]), enhancing the physical plausibility and artistic expressiveness for the generated dance. (2) Appearance Expert. Wan-Animate [5] has recently garnered substantial attention in both industry and academia. However, directly applying it to dance video generation yields limited effectiveness, as dance videos exhibit significantly more complex patterns than general videos. Thus, the Appearance Expert adopts a decoupled Kinematic–Aesthetic two-stage fine-tuning strategy to achieve high-fidelity appearance synthesis for dance videos. In Kinematic stage, it fine-tunes the Body Adapter to strengthen kinematic conditioning and motion adherence. In Aesthetic stage, it attaches a LoRA [11] branch to each DiT block and fine-tunes for aesthetic refinement, enhancing texture fidelity and stylistic consistency.
@@
To better benchmark music-driven dance video generation task, we curate a large-scale and diverse dataset and design a motion–appearance evaluation protocol. Firstly, we curate a large-scale dance video dataset, named MA-Data, comprising  70k clips of 5–10 seconds each (totaling 116 hours), spanning over 20 dance genres. The dataset consists of two complementary sources: (1) 3D-rendered data (motion-centric): Derived from FineDance [17]—the largest 3D dance dataset recorded by professional dancers—we render front-view videos and extract random 5–10 s segments via a sliding window, yielding  20k clips (28 h). This subset emphasizes dance motion rather than visual appearance. (2) In-the-wild internet data (appearance-centric): Collected from high-engagement videos on platforms such as TikTok and YouTube, using the same sliding-window strategy to obtain  50k 5-10 s clips (88 h). This subset emphasizes visual appearance, while motions are relatively unprofessional. Secondly, we introduce a motion–appearance evaluation protocol. For the motion dimension, we assess the fidelity, diversity, and synchronization [16, 17] from Human-Kinematics perspective based on the 2D keypoints extracted by ViTPose [41]. For the appearance dimension, we leverage VBench [14] to evaluate key visual factors critical to dance videos, including imaging quality, aesthetic quality, subject consistency, background consistency, motion smoothness, temporal flickering.
In conclusion, our contributions are as follows:
  • To better benchmark the music-driven dance video generation task, we curate a large-scale and diverse dataset named MA-Data, along with a motion–appearance evaluation protocol.
  • Based on them, we introduce MACE-Dance, a music-driven dance video generation framework with cascaded experts, achieving state-of-the-art (SOTA) performance.
  • The Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training strategy, achieving state-of-the-art (SOTA) performance in both quality and efficiency on the FineDance dataset in 3D dance generation task.
  • The Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance on the MA-Data dataset in the pose-driven image animation task.
2 Related Work
2.1 Music-Driven 3D Dance Generation
@@
Music and dance are deeply intertwined, and recent progress in music-to-dance generation has largely centered on 3D motion. Broadly, existing methods fall into three families: GAN-based, autoregressive, and diffusion-based models. 1) GAN-based models. Generators synthesize motion from music while discriminators provide adversarial feedback. Examples include CoheDancers [44], DeepDance [35], and Choreography cGAN [13]. 2) Autoregressive models. These methods typically adopt a two-stage pipeline: curating choreographic units by VQ-VAE [38] or FSQ [23], followed by autoregressive modeling of music-conditioned distributions over these units. Works such as Bailando [30], Bailando++ [31], Duolando [32], and MEGADance [47] fall into this paradigm. 3) Diffusion-based models. These methods corrupt motion with noise and train denoising networks to iteratively recover sequences conditioned on music, enabling diverse and temporally coherent dances. Representative works include EDGE [37], FineNet [17], Lodge [19], Lodge++ [18], and GCDance [20]. Despite substantial progress, 3D dance generation largely focuses on motion generation and underemphasizes visual appearance—an essential aspect of dance as an art form. Although 2D dance videos can be rendered from 3D motion, the outputs typically lack realistic human–scene interactions and high-fidelity human textures.
2.2 Human-Centric Image Animation
@@
In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance, making it a promising direction for dance video generation. Firstly, pose-driven image animation utilizes 2D keypoints to generate motion videos, achieving notable advances [36, 12, 5], including Animate-X [36], Animate Anyone [12] and Wan-Animate [5]. However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—remains manual [2]. Secondly, speech-driven image animation employs audio features to generate talking head videos, also achieving significant breakthroughs [26, 27, 28], such as SyncTalk [26], OmniSync [28] and Hallo2 [6]. However, its direct transfer to dance video generation remains challenging, as these methods primarily focus on relatively simple upper-body gestures rather than the complex full-body motion required in dance [25, 24]. Finally, research on music-driven dance video generation remains limited  [4, 40]. DabFusion [40] introduces an optical-flow–prediction-based Diffusion Model; however, the generated videos exhibit blurry foreground subjects and backgrounds, thereby degrading visual fidelity. X-Dancer  [4] first predicts 2D keypoints from music and then drives image animation with these keypoints. Although this yields modest improvements, it overlooks the inherently 3D nature of dance; relying on a 2D representation severely constrains the model’s understanding of dancer locomotion, ultimately compromising motion quality. In conclusion, existing works for dance video generation still fail to simultaneously deliver high-quality visual appearance and human motion. Thus, leveraging synergistic collaboration among cascaded experts, MACE-Dance generates dance videos that not only exhibit kinematically plausible and artistically expressive motion but also maintain spatiotemporally coherent appearance.
3 Methodology
3.1 Overview
@@
Given a music $M\in R^{T\times C_{m}}$ and reference image $I\in R^{H\times W\times 3}$, our objective is to synthesize the corresponding dance videos $D\in R^{T\times H\times W\times 3}$ with high-quality visual appearance and human motion. Overall, MACE-Dance is a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.  2. The Motion Expert (ME) transfers music sequence $M$ into 3D motion sequence $X\in R^{T\times C_{x}}$, enforcing kinematic plausibility and artistic expressiveness. The Appearance Expert (AE) utilizes the above 3D motion sequence $X$ and reference image $I$ to drive video synthesis, preserving visual identity with spatiotemporal coherence.
@@
Figure 2: Overview of MACE-Dance.Leveraging the cascaded Mixture-of-Experts (MoE) design, the Motion Expert generates kinematically plausible and artistically expressive 3D motion $X$ conditioned on the music $M$, and the Appearance Expert then animates the reference image II with the 3D motion $X$, yielding the dance video $D$ that exhibits spatiotemporally coherent appearance. Thanks to the Guidance-Free Training (GFT) strategy, $\beta\in[0,1]$ can serve as a controllable knob that governs the diversity of the generated motion.
3.2 Motion Expert
3.2.1 Generative Strategy.

### v3

@@ Abstract.
With the rise of online dance‑video platforms and rapid advances in AIGC, music‑driven dance generation task has emerged as a compelling research direction. Despite substantial progress in related domains such as music-driven 3D dance generation, pose-driven image animation, and audio-driven talking-head synthesis, these approaches are not readily transferable to this task due to fundamental mismatches in generation targets and constraints. Moreover, research on music-driven dance video generation remains limited and fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance. Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training (GFT) strategy, achieving state-of-the-art (SOTA) performance in 3D dance generation task; the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy, achieving state-of-the-art (SOTA) performance in the pose-driven image animation task. To better benchmark this task, we curate a large-scale dataset, and design a motion–appearance evaluation protocol. Based on them, MACE-Dance also achieves the state-of-the-art (SOTA) performance. Code is available at https://github.com/AMAP-ML/MACE-Dance.
copyright: noneccs: Applied computing Arts and humanitiesccs: Computing methodologies Computer visionccs: Human-centered computing
@@
Figure 1. Leveraging the synergistic collaboration among the cascaded experts, MACE-Dance can generate diverse dance videos that not only exhibit kinematically plausible and artistically expressive motion, but also maintain spatiotemporal coherent appearance.
1. Introduction
@@
Dance is a vital part of human culture. Moving to the beat and melody, dancers both convey emotion and narrative intent and showcase the power and beauty of human movement (Tseng et al., 2023; Butterworth*, 2004). In the era of the internet, dance videos have become highly prominent on platforms such as YouTube and TikTok. In parallel, rapid advances (Yang et al., 2025a; Zhuo et al., 2023; Chen et al., 2025a, b; Lei et al., 2025) in AI-generated content (AIGC) has created the technical preconditions for automating dance video creation, making it a timely and impactful research direction. Nevertheless, this task faces two key challenges: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency.
@@
Recent progress in dance generation has focused primarily on 3D dance (Tseng et al., 2023; Li et al., 2024c, 2023), with numerous strong methods emerging across model families-autoregressive (Siyao et al., 2022; Yang et al., 2025b, a), GAN-based (Yang et al., 2024b; Sun et al., 2019; Huang and Liu, 2021), and diffusion-based (Tseng et al., 2023; Li et al., 2024c, b). Although 2D dance videos can be rendered from 3D motion, such renderings typically lack realistic human–scene interactions and detailed appearance cues, resulting in visually suboptimal outputs (Yang et al., 2024c). In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos. Traditionally, pose-driven image animation has achieved notable advances (Tan et al., 2024; Hu, 2024; Cheng et al., 2025). However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step—still remains manual (Butterworth*, 2004). Similarly, audio-driven talking head generation has also achieved significant breakthroughs (Peng et al., 2024, 2025b, 2025c). However, its direct transfer to dance video generation remains challenging, as it primarily focuses on relatively simple upper-body gesture rather than the complex full-body motion required in dance (Peng et al., 2023). Research on music-driven dance video generation remains limited (Chen et al., 2025e; Wang et al., 2025b; Tang et al., 2025) and fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance.
@@
Accordingly, we present MACE-Dance, a music-driven dance video generation framework with cascaded mixture-of-experts (MoE), as shown in Fig.  1. The Motion Expert performs music-to-3D motion enforcing kinematic plausibility and artistic expressiveness, while the Appearance Expert carries out motion-and-reference conditioned video synthesis, preserving visual identity with spatiotemporal coherence. Notably, MACE-Dance adopts 3D SMPL (Loper et al., 2023) parameters rather than 2D keypoints as the intermediate representation, as 3D provides view-invariant and physically consistent supervision, while 2D projections introduce irreversible information loss and viewpoint ambiguity. (1) Motion Expert. Motion Expert adopts Diffusion Model with BiMamba-Transfomer hybrid architecture. The bidirectional Mamba (Gu and Dao, 2023) captures intra-modal local dependencies in music or dance, while the Transformer (Vaswani, 2017) models cross-modal global context. Owing to this architecture, the Motion Expert generates entire sequence in non-autoregressive manner during inference, not only improving generation efficiency, but also avoiding the exposure bias problem in autoregressive (Yang et al., 2025b) and inpainting-based (Tseng et al., 2023) methods. To enhance generation stability and accelerate inference, we employ guidance-free training (GFT (Chen et al., 2025c)) instead of conventional classifier-free guidance (CFG (Ho and Salimans, 2022)), enhancing the physical plausibility and artistic expressiveness for the generated dance. (2) Appearance Expert. Wan-Animate (Cheng et al., 2025) has recently garnered substantial attention in both industry and academia. However, directly applying it to dance video generation yields limited effectiveness, as dance videos exhibit significantly more complex patterns than general videos. Thus, the Appearance Expert adopts a decoupled Kinematic–Aesthetic two-stage fine-tuning strategy to achieve high-fidelity appearance synthesis. In Kinematic stage, it fine-tunes the Body Adapter to strengthen kinematic conditioning and motion adherence. In Aesthetic stage, it attaches a LoRA (Hu et al., 2022) branch to each DiT block and fine-tunes for aesthetic refinement, enhancing texture fidelity and stylistic consistency.
@@
To better benchmark music-driven dance video generation task, we curate a large-scale dataset and design a motion–appearance evaluation protocol. Firstly, we curate a large-scale dance video dataset, named MA-Data, comprising  70k clips of 5–10 seconds each (totaling 116 hours), spanning over 20 dance genres. The dataset consists of two complementary sources: (1) 3D-rendered data (motion-centric): Derived from FineDance (Li et al., 2023)—the largest 3D dance dataset recorded by professional dancers—we render front-view videos and extract random 5–10 s segments via a sliding window, yielding  20k clips (28 h). This subset emphasizes professional dance motion rather than visual appearance. (2) In-the-wild internet data (appearance-centric): Collected from high-engagement videos on platforms such as TikTok and YouTube, using the same sliding-window strategy to obtain  50k 5-10 s clips (88 h). This subset emphasizes visual appearance, while motions are relatively unprofessional. Secondly, we introduce a motion–appearance evaluation protocol. For the motion dimension, we assess the fidelity, diversity, and synchronization (Li et al., 2021, 2023) from Human-Kinematics perspective based on the 2D keypoints extracted by ViTPose (Xu et al., 2022). For the appearance dimension, we adopt VBench (Huang et al., 2024)—a widely used benchmark in video generation—and select a set of dance-specific metrics.
In conclusion, our contributions are as follows: (1) To better benchmark the music-driven dance video generation task, we curate a large-scale dataset named MA-Data, along with a motion–appearance evaluation protocol. (2) Based on them, we introduce MACE-Dance, a music-driven dance video generation framework with cascaded experts, achieving SOTA performance. (3) The Motion Expert adopts Diffusion Model with BiMamba-Transformer hybrid architecture and Guidance-Free Training strategy, achieving SOTA performance on the FineDance dataset in music-driven 3D dance generation task. (4) Appearance Expert adopts a decoupled Kinematic-Aesthetic fine-tuning strategy, achieving SOTA performance on the MA-Data dataset in the pose-driven image animation task.
2. Related Work
2.1. Music-Driven 3D Dance Generation
@@
Music and dance are deeply intertwined, and recent progress in music-to-dance generation has largely centered on 3D motion. Broadly, existing methods fall into three families: GAN-based, autoregressive, and diffusion-based models. 1) GAN-based models. Generators synthesize motion from music while discriminators provide adversarial feedback. Examples include CoheDancers (Yang et al., 2024b) and DeepDance (Sun et al., 2019). 2) Autoregressive models. These methods typically adopt a two-stage pipeline: curating choreographic units by VQ-VAE (van den Oord et al., 2017) or FSQ (Mentzer et al., 2023), followed by autoregressive modeling of music-conditioned distributions over these units (Yang et al., 2024a, 2026; Li et al., 2024a). Works such as Bailando (Siyao et al., 2022), Bailando++ (Siyao et al., 2023), and MEGADance (Yang et al., 2025b) fall into this paradigm. 3) Diffusion-based models. These methods corrupt motion with noise and train denoising networks to iteratively recover sequences conditioned on music (Yang et al., 2025c), enabling diverse and temporally coherent dances. Representative works include EDGE (Tseng et al., 2023), FineNet (Li et al., 2023), Lodge (Li et al., 2024c), Lodge++ (Li et al., 2024b), and GCDance (Liu et al., 2025). Despite substantial progress, 3D dance generation only focuses on motion generation and underemphasizes visual appearance—an essential aspect of dance as an art form. Although 2D dance videos can be rendered from 3D motion, the outputs typically lack realistic human–scene interactions and high-fidelity human textures.
2.2. Human-Centric Image Animation
@@
In contrast, human-centric image animation leverages a reference image along with various driving signals to generate videos that exhibit high-quality visual appearance, making it a promising direction for dance video generation. Firstly, pose-driven image animation utilizes 2D keypoints to generate motion videos, achieving notable advances (Tan et al., 2024; Hu, 2024; Cheng et al., 2025), including Animate-X (Tan et al., 2024), Animate Anyone (Hu, 2024) and Wan-Animate (Cheng et al., 2025). However, its utility for dance video generation is limited, as pose design—widely regarded as the most challenging and time-consuming step— still remains manual (Butterworth*, 2004). Secondly, speech-driven image animation employs audio features to generate talking head videos, also achieving significant breakthroughs (Peng et al., 2024, 2025b, 2025c), such as SyncTalk (Peng et al., 2024), OmniSync (Peng et al., 2025c) and Hallo2 (Cui et al., 2024). However, its direct transfer to dance video generation remains challenging, as these methods primarily focus on relatively simple upper-body gestures rather than the complex full-body motion required in dance (Peng et al., 2023; Zhang et al., 2025c, b, d). Finally, research on music-driven dance video generation remains limited. DabFusion (Wang et al., 2025b) introduces an end-to-end Diffusion-based method, but the generated videos exhibit blurry foreground subjects and backgrounds, thereby degrading visual fidelity. X-Dancer  (Chen et al., 2025e), STG-Mamba (Tang et al., 2025) and ChoreoMuse (Wang et al., 2025a) predict 2D keypoints from music and then drives image animation with these keypoints. However, they remain limited in handling limb occlusions and complex full-body locomotion in dance videos. In conclusion, existing works for dance video generation still fails to capture the inherently 3D nature of dance, resulting in compromised motion quality and visual appearance. Thus, we propose MACE-Dance, a cascaded expert framework that synergistically integrates motion and appearance generation, producing kinematically plausible and artistically expressive motion while maintaining spatiotemporally coherent visual appearance.
3. Methodology
3.1. Overview
@@
Given a music $M\in R^{T\times C_{m}}$ and reference image $I\in R^{H\times W\times 3}$, our objective is to synthesize the corresponding dance videos $D\in R^{T\times H\times W\times 3}$ with high-quality visual appearance and human motion. Overall, MACE-Dance is with cascaded mixture-of-experts (MoE), as shown in Fig.  2. The Motion Expert (ME) transfers music sequence $M$ into 3D motion sequence $X\in R^{T\times C_{x}}$, enforcing kinematic plausibility and artistic expressiveness. The Appearance Expert (AE) utilizes the above 3D motion sequence $X$ and reference image $I$ to drive video synthesis, preserving visual identity with spatiotemporal coherence. This task decoupling significantly reduces the complexity of learning a direct music-to-video mapping by isolating motion semantics from visual appearance. Moreover, the explicit 3D motion representation suppresses spurious cross-modal correlations and provides an interpretable intermediate interface for robust and controllable video synthesis.
Unlike prior works (Chen et al., 2025e; Tang et al., 2025) that adopt 2D keypoints as the intermediate representation, we instead use 3D motion as the bridge between the two experts for three reasons. (1) Richer spatial fidelity. 3D motion preserves full-body geometric structure, including global translation and orientation, which is essential for dance phrases with large-amplitude locomotion and complex spatial choreography, whereas 2D projections inevitably discard depth and global movement information. (2) Cleaner supervision. 3D representation disentangles pose from camera viewpoint and subject-specific appearance, providing a more stable and generalizable signal for learning the music-to-motion correspondence, while 2D keypoints are entangled with perspective and body proportions. (3) Better robustness. 3D motion is inherently more robust to self-occlusion and viewpoint variation, whereas 2D poses often suffer from missing joints, depth ambiguity, and inconsistent observations. Additionally, we adopt SMPL (Loper et al., 2023) as the representation of the 3D motion sequence $X$ for two reasons. (1) Prior focus on body motion. Most existing 3D dance generation methods primarily model body-level motion rather than detailed hand articulation. In our setting, body-level motion alone is sufficient to produce strong visual results, as also evidenced by our demo videos. (2) Extensibility. Our framework can be readily extended to richer motion representations, such as SMPL-X, when suitable data become available.
@@
Figure 2. Overview of MACE-Dance.Leveraging the cascaded Mixture-of-Experts (MoE) design, the Motion Expert generates kinematically plausible and artistically expressive 3D motion $X$ conditioned on the music $M$, and the Appearance Expert then animates the reference image II with the 3D motion $X$, yielding the dance video $D$ that exhibits spatiotemporally coherent appearance. Thanks to the Guidance-Free Training (GFT) strategy, $\beta\in[0,1]$ can serve as a controllable knob that governs the diversity of the generated motion.
3.2. Motion Expert
3.2.1. Generative Strategy.
Diffusion.